Physical Units Method

Alomar Company manufactures four products from a joint production process: barlon, selene, plicene, and corsol. The joint costs for one batch are as follows:

Direct materials $73,659
Direct labor 37,301
Overhead 22,246

At the split-off point, a batch yields 1,589 barlon, 4,189 selene, 3,467 plicene, and 5,201 corsol. All products are sold at the split-off point: barlon sells for $17 per unit, selene sells for $20 per unit, plicene sells for $27 per unit, and corsol sells for $33 per unit.

Required:

1. Allocate the joint costs using the physical units method. If required, round your percentage allocation to four decimal places and round allocated costs to the nearest dollar. Note: The total of the allocated cost does not equal to the one provided in the question data due to rounding error.

Allocated Joint Cost
Barlon $
Selene
Plicene
Corsol
Total $

2. Suppose that the products are weighted as shown below:

Barlon 1.0
Selene 2.0
Plicene 1.6
Corsol 2.4

Allocate the joint costs using the weighted average method. If required, round your percentage allocation to four decimal places and round allocated costs to the nearest dollar.

Allocated Joint Cost
Barlon $
Selene
Plicene
Corsol
Total $
Answer #1

(1) physical units method allocates cost to products in proportion to the units of product produced.

Total units = 1,589+4,189+3,467+5,201

=14,446

Joint cost = $73,659+$37,301+$22,246

=$133,206

units percent (units/total units) Allocated costs(Total joint cost*percent)
Barlon 1589 0.1100(1589/14,446) $14,653(0.1100*$133,206)
Selene 4189 0.2900(4189/14,446) $38,630(0.2900*$133,206)
Plicene 3467 0.2400(3467/14,446) $31,969(0.2400*$133,206)
Corsol 5201 0.3600(5201/14446) $47,954(0.3600*$133,206)
Total 14,446 $133,206

2) physical units are applied to the obtain percentage of weighted cases.percentage are multiplied by join costs to avail allocated joint costs.

Total weight = 1+2+1.6+2.4=7.0

Weight Percent Allocated Joint Cost
Barlon 1.0 0.1429(1/7) $19,035 ($133,206*0.1429)
Selene 2.0 0.2857(2/7) $38,057 ($133,206*0.2857)
Plicene 1.6 0.2286 (1.6/7) $30,451 ($133,206*0.2286)
Corsol 2.4 0.3429(2.4/7) $45,676 ($133,206*0.3429)
Total 7.0 $133,219